Two trains A and B are moving towards each other at a speed of 156.2m/s. If the frequency of the whistle emitted by A is 844.6Hz, then what is the apparent frequency of the whistle heard by the passenger sitting in train B. Note that the velocity of sound in air is 340m/s.


1
Expert's answer
2021-12-08T11:30:34-0500

"\nu\ '=\nu_0\frac{v+v\ '}{v-u}=844.6\cdot\frac{340+156.2}{340-156.2}=2280\ (Hz)" . Answer